it takes julie 60 minutes to pick the staples out of the wall in the hallway. Zach can do the same job in 40 minutes . how long will it take them if they work together?

Answer by richwmiller(17219)

1/60+1/40=1/x
2/120+3/120=1/x
5/120=1/x
120/5=x
24=x